OpenStudy (anonymous):

For homework it says simplify 5-3(4-3tothe2ndpower) And the answer is 20, but I can't figure out the steps to take to get 20

OpenStudy (anonymous):

remember PEMDAS parentheses, exponents, multiplication and division, addition and subtraction so written out: \[5-3(4-3^2) = 5-3(4-9)=5-3(-5)=5-(-15)=5+15=20\]